# On the Foundation of Distributionally Robust Reinforcement Learning

### v3

@@ Abstract
Motivated by the need for a robust policy in the face of environment shifts between training and deployment, we contribute to the theoretical foundation of distributionally robust reinforcement learning (DRRL). This is accomplished through a comprehensive modeling framework centered around distributionally robust Markov decision processes (DRMDPs). This framework obliges the decision maker to choose an optimal policy under the worst-case distributional shift orchestrated by an adversary. By unifying and extending existing formulations, we rigorously construct DRMDPs that embrace various modeling attributes for both the decision maker and the adversary. These attributes include adaptability granularity, exploring history-dependent, Markov, and Markov time-homogeneous decision maker and adversary dynamics. Additionally, we delve into the flexibility of shifts induced by the adversary, examining SA and S-rectangularity. Within this DRMDP framework, we investigate conditions for the existence or absence of the dynamic programming principle (DPP). From an algorithmic standpoint, the existence of DPP holds significant implications, as the vast majority of existing data and computationally efficient RL algorithms are reliant on the DPP. To study its existence, we comprehensively examine combinations of controller and adversary attributes, providing streamlined proofs grounded in a unified methodology. We also offer counterexamples for settings in which a DPP with full generality is absent.
@@ 1 Introduction
Our current era has growingly witnessed the benefits of data-driven sequential decision making, which has unleashed – and has much further potential to unleash – productivity forces in a diverse range of applications of societal impact. Such applications range from the traditional engineering domains such as automation (Landgraf et al., 2021), inventory management (Gong and Simchi-Levi, 2023; Mao et al., 2020), and control of service systems (Hu et al., 2021; Zhalechian et al., 2023) to the newly emerged areas in the digital economy such as personalized product recommendation in online platforms (Afsar et al., 2022), feature-based dynamic pricing (Bertsimas and Perakis, 2006), and real-time bidding in high-frequency auctions (Grislain et al., 2019; Zhao et al., 2018).
In modern applications, the reliability of the learning process is crucial to ensure that a policy learned from historical data or simulations performs effectively when deployed in real-world scenarios. Classical reinforcement learning (RL), where the learning process is typically modeled as an unknown Markov decision process (MDP), has demonstrated remarkable empirical success in simulated environments, with application domains spanning robotics (Kormushev et al., 2013), autonomous driving (Shalev-Shwartz et al., 2016; Pan et al., 2017; Kiran et al., 2021), and games (Mnih et al., 2015; Silver et al., 2017, 2018). However, extending the success observed in simulated environments to applications in real environments poses challenges and remains an active area of research. A major obstacle is the frequently misplaced assumption that the training and deployment environments of RL policies are congruent. The reality, however, is rife with discrepancies caused by model mis-specifications, environment shifts, and various other confounding factors. Another challenge emerges from the intricacies of computing optimal policies in models where the complete Markov state of the system is either not fully observable to the modeler and, consequently, to the controller, or intentionally omitted due to computational considerations. To address these challenges, it is natural to explore distributionally robust (DR) formulations that take into account the impact on performance arising from the differences between the training and real environments.
This paper studies foundational aspects of distributionally robust reinforcement learning (DRRL), an approach designed to bridge the chasm between idealistic models and the complex reality. At the heart of our formulation lies a strategic game. One agent, representing the controller, desires to maximize the expected cumulative reward. In contrast, the other agent, severing as the adversary that embodies potential environmental shifts, selects the stochastic environment with the aim of diminishing this reward. This adversarial interaction model proves insightful in capturing the complexities associated with designing a reliable controller for real environments, especially when relying on a simplified training environment. Hence, DRRL provides a more resilient and realistic training paradigm, addressing the challenges that arise in deploying policies to real-world application.
A key contribution of our research lies in presenting a unified approach to both controller and adversary models, incorporating salient attributes. This broadens the spectrum of behaviors encapsulated by our formulations of DRMDP, and hence enhances their applicability. Moreover, our investigation delves into various facets of the interactions between the controller and the adversary. One aspect is the information availability to the controller and the adversary at the time of action, encompassing considerations of history-dependence, Markov, and time-homogeneity. For instance, within a simulated MDP environment, the use of time-homogeneous adversary implies that the controllers trained in the simulator assumes the real environment is also an MDP with the same states and actions, although likely with different and unknown transition dynamics. In this setting, the corresponding DRRL problem aims to learn from the simulator a best controller policy for the worst shift in a predetermined set of possible environments.
In certain applications, there is a legitimate concern that the true complexity of the real-world environment far exceeds the simplicity of the Markov model, particularly with respect to the states considered. For example, a simulated MDP environment may fail to capture some hidden states of the real environment, making it not an MDP for the controller when using the same state information as in the simulator. This learning settings necessitates the consideration of a richer model of how the distributional shift occurs. As such, even if the modeler doesn’t posit the presence of an actual adversary actively harming the controller, the inclusion of a history-dependent controller and adversary model is beneficial: this model accommodates the non-Markov nature of the partially observed system, which comprises latent states inaccessible to the controller but influential in driving the system’s evolution. For further elaborations, see the proceeding Section 1.2. This addresses a distinct source of the simulator-to-real-environment gap, one that cannot be effectively tackled by the time-homogeneous adversary mentioned earlier.
Additionally, we examine attributes of the adversary necessary for an accurate DRRL model, aiming to strike a balance between capturing the essential features of the real environment and maintaining computational feasibility. Concretely, an effective adversary in this context must navigate the challenge of addressing environment shifts while preserving the structural integrity of real environments and avoiding overly conservative policies. Striking the right balance is crucial: an adversary that is too weak may fail to encompass the full spectrum of environments an RL agent might encounter, thereby undermining the model’s reliability in the deployment environment. On the other hand, an adversary that is too powerful might compel the RL agent to consider an array of improbable scenarios, leading to the violation of the structural integrity of the modeled system and the adoption of overly cautious and impractical policies. Our framework provides the flexibility to tailor an adversary to the problem by integrating various specific attributes, such as rectangular adversarial decisions and the imposition of convexity in the adversarial action set.
In this paper, our central objective is to comprehensively investigate the conditions under which formulations of DRRL, accounting for the aforementioned variety of attributes for both the controller and the adversary, result in a dynamic programming principle (DPP). This is expressed in the form of a Bellman-type equation (Bellman, 1954). This equation, once established, will serve as a foundation for the subsequent development and analysis of efficient RL algorithms—a domain we plan to explore in subsequent research endeavors.
This paper achieves this goal by establishing a unifying framework for the verification of the existence or absence of a DPP. A more detailed exposition of our results is presented in the upcoming discussion. Moreover, our framework provides guidelines for both theorists and practitioners in the selection and design of DRRL environments, encompassing a diverse spectrum of controller and adversary attributes, as presented in Section 1.2. The flexibility within our framework facilitates the development of practical DRRL algorithms, both model-based and model-free, while simultaneously yielding policies that effectively hedge against the unpredictability inherent in real-world applications.
@@ 1.1 Results and Methodology
We first present an overview of the fundamental concepts in this manuscript. The study addresses the problem of distributionally robust Markov decision processes (DRMDP) with infinite-horizon $\gamma$-discounted reward, formulated as follows:
@@
where $r$ is the reward function, $X_{t}$ and $A_{t}$ represent the state and action at time $t$ respectively, and $X_{0}\sim\mu$ as the initial distribution. The controller selects a policy $\pi$ from the class $\Pi$, maximizing the worst-case reward against an adversary who, with knowledge of the controller’s policy, selects a policy from a class $\mathrm{K}$ Here, the policy classes $(\Pi,\mathrm{K})$ considered in this paper will be formally defined in Section 2. Despite extensive exploration in the literature (Nilim and El Ghaoui, 2005; González-Trejo et al., 2002; Iyengar, 2005; Xu and Mannor, 2010; Le Tallec, 2007; Wiesemann et al., 2013; Shapiro, 2021; Goyal and Grand-Clément, 2023; Li and Shapiro, 2023), there remains ambiguity in the rigorous formulation of Problem 1.1. This paper contributes to this literature by defining and examining the dynamic behavior arising from various combinations of controllers and adversaries. The investigation takes into account variations in information availability and rectangularity, shedding light on nuanced aspects that enrich our understanding of the DRMDP problem.
In terms of information perspectives, we systematically examine history-dependent, Markov, and Markov time-homogeneous controllers and adversaries. A history-dependent entity in either role has the flexibility to choose different actions or distributions based on its historical information. In contrast, a Markov controller or adversary lacks memory of prior states and actions but retains the capacity to take time-and-state-dependent actions or distributions. Further narrowing the scope, a Markov time-homogeneous controller or adversary relinquishes information about time, compelling it to consistently employ a static state-dependent action or distribution, starting from time 0 and persisting thereafter.
The notion of rectangularity, as initially introduced in the DRMDP literature, originally pertained to the adversary’s freedom to vary actions over time (Iyengar, 2005). However, owing to the adoption of the information perspectives we discussed earlier and the formulation characteristics allowing for the strategic use of rectangularity to limit the adversary’s power, the concept has evolved into a mechanism for imposing structural constraints on the adversary. This conceptual change is articulated in Le Tallec (2007) and Wiesemann et al. (2013). In the DRMDP literature, the prevalent assumptions are S and SA-rectangularity. An SA-rectangular adversary enjoys the liberty to choose different action distributions for each current state-action pair $(s,a)\in S\times A$. On the contrary, S-rectangularity restricts this freedom by permitting the adversary to choose its action distribution with prescribed dependence among each action $a\in A$ within the same state $s$. Example 1 provides a demonstration of the difference between an SA and an S-rectangular adversary. Additionally, it is conceivable to further limit the adversary’s power by mandating that their choices maintain predefined dependences among all state and action pairs. This type of structural constraint is termed general-rectangularity.
The preceding explanations offer heuristic insights into the pivotal concepts recurrently discussed in this paper. For precise and formal definitions, please refer to Section 2.
@@
As mentioned earlier, the focus of this paper is to identify scenarios in which the optimal policy satisfies the distributionally robust variant DPP, a.k.a. distributionally robust Bellman equation:
$u(s)=\sup_{d\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}E_{a\sim d}[r(s,a)]+\gamma E_{s^{\prime}\sim d\otimes p_{s}}[u(s^{\prime})],\quad s\in S,$ (1.2)
@@
where $\otimes$ denotes the product of measures and $\mathcal{Q}$ represents the set of probability distributions if the controller is allowed to be randomized, whereas $\mathcal{Q}$ denotes the action set (understood as Dirac measures) when the controller is limited to being deterministic. $\mathcal{P}_{s}$ is the adversary’s action distribution set (also known as an ambiguity set in the DRMDP literature) at state $s$.
@@
In view of this objective, it is natural to assume an SA or S-rectangular adversary, as achieving a DPP with complete generality might not be feasible when dealing with a general-rectangular adversary (Le Tallec, 2007; Wiesemann et al., 2013).
@@
      Controller History- González-Trejo et al. (2002) Iyengar (2005)
@@
    Markov Nilim and El Ghaoui (2005) Nilim and El Ghaoui (2005)
    Time- Iyengar (2005) Nilim and El Ghaoui (2005)
    homogeneous Nilim and El Ghaoui (2005)
@@
      Randomized Controller History- Xu and Mannor (2010) Wiesemann et al. (2013)
    dependent Xu and Mannor (2010)
    Markov Li and Shapiro (2023)
    Time- Le Tallec (2007)
@@
    Table 3: Randomized controller versus possibly non-convex S-rectangular adversary policy classes.
@@
      Randomized Controller History- Wiesemann et al. (2013)
@@
    Markov Li and Shapiro (2023)
    Time- Le Tallec (2007)
@@
  • Table 4 also assumes an S-rectangular taking actions in a convex or possibly non-convex set. However, the controller is constrained to choose deterministic actions.
@@
As summarized in Section 3.4, the unifying principles stem from two perspectives: the convexity of the action sets of the controller and the adversary, and the interchangeability of the sup-inf within the DR Bellman equation (3.1).
@@
From the convexity viewpoint, the six cases located in the lower triangles of all tables consistently hold without requiring convexity assumptions, as proved in Theorem 1. If the controller opts for randomized actions from a convex set of distributions, then, additionally, the scenario where a history-dependent controller interacts with an S-rectangular Markov adversary with a possibly non-convex action set results in a DPP. It’s important to note that even if the adversary’s action set is convex, a non-convex controller could lead to the same non-existence of DPPs as in the non-convex adversary case. This is also indicated in Table 4.
From the interchangeability perspective, if the interchanged Bellman equation (3.2) yields the same solution as (3.1), then the DPPs hold for all 9 cases of information structures of the two parties. This includes the case where the adversary is SA-rectangular, validating Table 1.
@@
The convergence of these two perspectives occurs when the conditions of Sion’s min-max principle (Sion, 1958) are met—specifically, when both parties have convex action distribution sets, and one of them is compact. Consequently, the interchanged equation (3.2) has the same solution as (3.1), ensuring that all 9 cases adhere to their respective DPPs. This results in the validity of Table 2.
@@
We further establish the anticipated implication of the DPP: the optimality of a Markov time-homogeneous controller. Moreover, we demonstrate that such an optimal policy can be derived directly from the solution to the DR Bellman equation. Consequently, we establish that the satisfaction of DPP validates the DRRL approaches that aim to learn an optimal robust policy by approximately solving the DR Bellman equation, either for the value function (3.1) in the S-rectangular case, or for the $q$-function (3.4) in the SA-rectangular case.

### v4

@@ Abstract
Motivated by the need for a robust policy in the face of environment shifts between training and deployment, we contribute to the theoretical foundation of distributionally robust reinforcement learning (DRRL). This is accomplished through a comprehensive modeling framework centered around robust Markov decision processes (RMDPs). This framework obliges the decision maker to choose an optimal policy under the worst-case distributional shift orchestrated by an adversary. By unifying and extending existing formulations, we rigorously construct RMDPs that embrace various modeling attributes for both the decision maker and the adversary. These attributes include the structure of information availability—covering history-dependent, Markov, and Markov time-homogeneous dynamics—as well as constraints on the shifts induced by the adversary, with a focus on SA- and S-rectangularity. Within this RMDP framework, we investigate conditions for the existence or absence of the dynamic programming principle (DPP). From an algorithmic standpoint, the existence of DPP holds significant implications, as the vast majority of existing data and computationally efficient DRRL algorithms are reliant on the DPP. To investigate its existence, we systematically analyze various combinations of controller and adversary attributes, presenting streamlined proofs based on a unified methodology. We then construct counterexamples for settings where a fully general DPP fails to hold and establish asymptotically optimal history-dependent policies for key scenarios where the DPP is absent.
@@ 1 Introduction
Our era has increasingly witnessed the transformative impact of reinforcement learning (RL), which has the potential to deliver – and in some cases has already delivered – significant productivity gains across a wide range of applications. Progress and applications of RL span traditional engineering domains such as automation [32], inventory management [17, 42], and the control of service systems [26, 85], as well as emerging areas in the digital economy, including personalized product recommendations on online platforms [1], feature-based dynamic pricing [4], and real-time bidding in high-frequency auctions [21, 86].
For deployment of RL in real-world applications, the reliability of the learning process is crucial, i.e., the policy learned from historical data or simulations needs to perform effectively when deployed in real-world scenarios. Standard reinforcement learning (RL), where the environment is typically modeled as an unknown Markov decision process (MDP), has demonstrated remarkable empirical success in simulated environments, with application domains spanning robotics [31], autonomous driving [58, 50, 30], and games [44, 68, 67].
However, extending the success of RL in simulated environments to real environments remains challenging, largely due to the simplistic assumption, on which existing RL techniques critically depend, that the training and deployment environments are the same. In reality, it is common to encounter discrepancies arising from model mis-specifications (sim2real gap), environment shifts, and other confounding factors in real systems that could make the underlying dynamics non-Markovian are not captured by the simulator.
To address these challenges, we explore robust MDP (RMDP) formulations as a foundation for distributionally robust reinforcement learning (DRRL). We present a unified RMDP framework designed to support reliable policy learning under a wide range of model misspecifications, providing a principled approach to bridging the gap between idealized models and real-world complexities. At the core of our formulation is a two-player game: the controller seeks to maximize the expected cumulative rewards, while the adversary, modeling the possible yet unknown environmental shifts, selects a stochastic environment to hinder performance. This adversarial perspective offers a powerful lens for designing reliable controllers under uncertainty, particularly when the training occurs in potentially mis-specified environments (unknown unknowns), thereby promoting a more resilient learning paradigm.
Despite a flourishing literature, see Section 1.2 for a detailed discussion, on this topic, practitioners hoping to design a DRRL algorithm for their specific use case would be quickly and justifiably lost. There are two main reasons for this.
  1. 1.
    Modeling difficulty: Modeling the controller and the adversary, particularly the information each party has at the time of action, requires knowing – and is hence reflecting the assumptions of – the underlying application at hand. For example, modeling the adversary as time-homogeneous reflects the assumption or knowledge that the real environment is also an MDP, one that shares the same state and action spaces as the simulator, albeit with different and unknown transition probabilities. Here, the DRRL problem aims to learn a controller policy from the simulator that performs the best against the worst-case environment shift within a specified ambiguity set.
    However, in other applications, the practitioner might know beforehand that the real environment cannot be Markovian, as a result of the latent variables that are neither observed in the real system nor captured by the simulator. This calls for a richer model of the distributional shift, where incorporating a history-dependent adversary model can be valuable. Since the real environment is not fully observable, from the controller’s perspective, the environment’s dynamics will appear history-dependent. As such, positing that the environment is a history-dependent adversary allows the controller to robustify against this more complex environment. For details, see Section 8.
    Unfortunately, the existing literature lacks a clear exposition on this front. Often, a particular formulation is adopted throughout the paper without much discussion on the modeling aspect that led to it or the alternative formulations that one should consider under other modeling assumptions. Another important modeling feature that tends to be left out of the discussion by the existing literature is the ambiguity set/adversary’s decision set, which encodes how much power to grant to the adversary.
    An effective adversary must account for meaningful environment shifts without compromising the structural realism of the system or inducing overly conservative policies. Striking this balance is critical: a weak adversary may overlook relevant variations the RL agent could face in deployment, reducing robustness, whereas an overly strong adversary may force the agent to prepare for unrealistic worst-case scenarios, leading to impractical and overly cautious behavior.
  2. 2.
    Algorithm design difficulty: When the practitioner has thought through all the modeling aspects and finalized a particular RMDP model, how can he/she design a DRRL algorithm? Given that the dynamic programming principle (DPP) is the algorithmic foundation for any RL algorithm, one would naturally look into the corresponding DPP for the RMDP settings as the very first step. Unfortunately, the situation for RMDP models is more complex: not all formulations of RMDP that can reflect and represent the diverse modeling alternatives discussed above enjoy the DPP. This raises an important question for the practitioner: Which formulation satisfies the DPP, thereby making it tractable to design an effective DRRL algorithm, and which does not?
    Unfortunately, the existing literature does not provide definitive answers to this question, which has important consequences regardless of the answer being a yes or a no for a particular formulation the practitioner has at hand. If the answer is affirmative, a corresponding DRRL algorithm can likely be developed straightforwardly by, for example, converting the equality into an update rule. Conversely, even a negative answer is valuable. Just as knowing an optimization problem is nonconvex informs one to relax some constraints to achieve convexity, understanding the limitations of a particular formulation can guide adjustments in modeling assumptions toward one that satisfies the DPP. Such an iterative modeling approach, however, requires a comprehensive mapping of the DPP for DRRL, which is currently lacking.
Our paper aims to address these two types of difficulties and hence make the methodology of DRRL – which is sound on a philosophical level – practically relevant and prescriptive to a practitioner looking to build a robust controller. First, on the modelling side, a key contribution of our work is the development of a unified framework that systematically incorporates salient attributes of both the controller and the adversary, thereby expanding the range of behaviors captured by our RMDP formulations and enhancing their practical relevance. Our analysis explores multiple dimensions of the controller–adversary interaction, one of which is the availability of information to each party at the time of action. This includes considerations of history-dependent, Markov, and time-homogeneous structures. Additionally, we investigate the design of adversarial models in DRRL, focusing on how to capture essential features of real-world environments while preserving computational tractability. Our framework offers flexibility to navigate this trade-off by allowing the incorporation of specific structural properties, such as rectangularity and convexity in the adversarial decision sets, tailored to the application at hand.
Second, on the algorithm design side, we investigate the conditions under which DRRL formulations, accounting for the aforementioned variety of attributes for the controller-adversary dynamics, result in a dynamic programming principle (DPP) expressed through a Bellman-type equation [3]. As already mentioned, establishing the DPP provides the foundation for developing and analyzing efficient RL algorithms, even before any DRRL algorithm is actually designed. To this end, we introduce a unifying framework for verifying the existence or absence of a DPP across various settings. A detailed overview of our results follows in the next sections. Beyond theoretical insights, our framework offers practical guidance for both researchers and practitioners in selecting and designing DRRL environments, as detailed in Section 8. This flexibility supports the design of both model-based and model-free DRRL algorithms, enabling the construction of policies that effectively hedge against the uncertainties of real-world deployment.
@@ 1.1 Results and Methodology
We first present an overview of the fundamental concepts in this manuscript. The study addresses the problem optimal control of RMDPs with infinite-horizon $\gamma$-discounted max-min value, formulated as follows:
@@
where $r$ is the reward function, $X_{t}$ and $A_{t}$ represent the state and action at time $t$ respectively, and $X_{0}\sim\mu$ as the initial distribution. The controller selects a policy $\pi$ from the class $\Pi$, maximizing the worst-case reward against an adversary who, with knowledge of the controller’s policy, selects a policy from a class $\mathrm{K}$. Here, the policy classes $(\Pi,\mathrm{K})$ considered in this paper will be formally defined in Section 2.
Despite extensive exploration in the literature [47, 18, 28, 80, 34, 79, 59, 19, 37], a unified framework that captures the full range of formulations for Problem (1.1) is still lacking. This gap makes it difficult to understand the modeling implications of DRRL and to design algorithms that reliably attain optimal policies. Our paper addresses this need by systematically examining the dynamic behavior resulting from different controller-adversary combinations, considering variations in information availability and rectangularity, thereby enriching our understanding of the RMDP problem and facilitating the deployment of DRRL.
From an information perspective, we systematically construct and analyze classes of controller and adversary policies that are history-dependent, Markov, or Markov time-homogeneous. A history-dependent agent bases actions on past observations; a Markov agent uses only the current state and time; and a Markov time-homogeneous agent relies solely on the current state, applying the same rule across all time periods. A key innovation of our framework is that these information structures can be assigned asymmetrically—for example, a history-dependent controller facing a time-homogeneous adversary—enabling a rich family of robust decision-making models.
The concept of rectangularity, originally introduced in the RMDP literature to describe the adversary’s ability to vary actions over time [28], has since evolved. With the adoption of different information structures and a growing emphasis on limiting adversarial power, rectangularity is now used to impose structural constraints on the adversary, as discussed in Le Tallec, [34] and Wiesemann et al., [79]. Common assumptions in the literature include S-rectangularity and SA-rectangularity. An SA-rectangular adversary can choose distinct action distributions for each state-action pair $(s,a)\in S\times A$, while S-rectangularity imposes restrictions by requiring a consistent distribution across actions within each state. Example 1 illustrates this distinction.
@@
As mentioned earlier, the focus of this paper is to identify scenarios in which the max-min control value (1.1) satisfies a robust DPP, a.k.a. robust Bellman equation:
$u(s)=\sup_{\phi\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}[r(s,A_{0})+\gamma u(X_{1})],\quad s\in S,$ (1.2)
@@
where the distribution of $A_{0},X_{1}$ under $P_{\phi,p_{s}}$ is given by $P_{\phi,p_{s}}(A_{0}=a_{0},X_{1}=s_{1})=\phi(a_{0})p_{s,a_{0}}(s_{1})$. Further, if the action of the controller is allowed to be randomized, then $\mathcal{Q}$ represents the set of probability distributions, whereas $\mathcal{Q}$ will denote the action set (understood as Dirac measures) when the controller is limited to being deterministic. $\mathcal{P}_{s}$ is the adversary’s action distribution set (also known as an ambiguity set in the RMDP literature) at state $s$.
@@
In view of this objective, it is natural to assume an SA or S-rectangular adversary, as achieving a DPP with complete generality might not be feasible when dealing with a general-rectangular adversary [34, 79].
@@
         Controller History- González-Trejo et al., [18] Iyengar, [28]
@@
    Markov Nilim and El Ghaoui, [47] Nilim and El Ghaoui, [47]
    Time- Iyengar, [28] Nilim and El Ghaoui, [47]
    homogeneous Nilim and El Ghaoui, [47]
@@
         Randomized Controller History- Xu and Mannor, [80] Wiesemann et al., [79]
    dependent Xu and Mannor, [80]
    Markov Li and Shapiro, [37]
    Time- Le Tallec, [34]
@@
    Table 3: Randomized controller versus non-convex S-rectangular adversary policy classes.
@@
         Randomized Controller History- Wiesemann et al., [79]
@@
    Markov Li and Shapiro, [37]
    Time- Le Tallec, [34]
@@
  • Table 4 also assumes an S-rectangular adversary taking actions in a convex or possibly non-convex set. However, the controller is constrained to choose deterministic actions.
@@
As summarized in Section 3.4, the unifying principles stem from two perspectives: the convexity of the action sets of the controller and the adversary, and the interchangeability of the sup-inf within the robust Bellman equation (3.1).
@@
Convexity perspectives: The six cases located in the lower triangles of all tables consistently hold without requiring convexity assumptions, as proved in Theorem 1. If the controller opts for randomized actions from a convex set of distributions, then, additionally, the scenario where a history-dependent controller interacts with an S-rectangular Markov adversary with a possibly non-convex action set results in a DPP (c.f. Theorem 4). It’s important to note that even if the adversary’s action set is convex, a non-convex controller could lead to the same non-existence of DPPs as in the non-convex adversary case.
Interchangeability perspective: Under our unifying framework, we establish an important novel result in Theorem 2: if the interchanged Bellman equation (3.2) admits the same solution as (3.1), then the DPP holds universally, regardless of the information asymmetry between the controller and the adversary. This insight, in particular, confirms the validity of the DPP in settings where the adversary is SA-rectangular (Theorem 3), leading to the results in Table 1.
@@
The convergence of these two perspectives occurs when the conditions of Sion’s min-max principle [70] are met–specifically, when both parties have convex action distribution sets, and one of them is compact. Consequently, the interchanged equation (3.2) has the same solution as (3.1), ensuring that all 9 cases adhere to the DPP (Corollary 2.1). This results in the validity of Table 2.
@@
We further show that the DPP implies the optimality of a Markov time-homogeneous controller and that such a policy can be directly obtained from the solution to the robust Bellman equation. This confirms the validity of DRRL approaches that learn a robust optimal policy by approximately solving the robust Bellman equation—either the value function in the S-rectangular case (3.1) or the $q$-function in the SA-rectangular case (3.4).